# A stochastic optimization algorithm for revenue maximization in a service system with balking customers

### v1

@@ 3.2 Coupling of systems with different initial workloads
    $<\left(\overline{W}^{1}_{j}-\overline{W}^{2}_{j}\right).$ (3.5)
@@
    $N_{2}\left(0,t\right]-N_{1}\left(0,t\right]=N_{2}\left(t-\Delta,t\right]\leq_{\text{st}}\overline{N}\left(\Delta\right),$ (3.6)
    where the random variable $\overline{N}(\Delta)$ records the number of arrivals in an interval of length $\Delta$, under a Poisson stream of arrivals with rate $\Lambda H\big(\underline{p},0\big)$. The stochastic dominance follows from an evident coupling argument: each interarrival time in the second queue stochastically dominates an $\text{Exp}\big(\Lambda\,H\big(\underline{p},0\big)\big)$ random variable. Similarly, under the complementary event $\big\{\overline{W}^{1}_{0}<\overline{W}^{2}_{0}\big\}$ we have (3.6) as well. In conclusion, $\big|N_{1}\left(0,T^{*}\right]-N_{2}\left(0,T^{*}\right]\big|\leq_{\text{st}}\overline{N}\left(\Delta\right).$
@@
The coupling featuring in Proposition 3.9 is illustrated by Figure 5.
4 Gradient estimation
Recalling the gradient estimator (2.8), the main objectives of this section are (i) to show how its components are computed and (ii) to establish almost convergence of the estimator when the sampling window increases.
For $k\geq 1$, let $N_{k}$ be the random variable which denotes the number of effective arrivals during the $k$-th iteration, i.e., during the time interval $\big[\overline{T}_{k-1},\overline{T}_{k}\big)$, and let $\widetilde{N}_{k}=N_{1}+\cdots+N_{k}$. Let $\left\{T^{*}_{k}\right\}_{k\geq 1}\in\mathbb{R}_{+}$ be a strictly increasing deterministic sequence, chosen by the service operator. Since $T_{k}$ is a random variable, the service operator cannot choose its value. Therefore, a quantity $T^{*}_{k}$ is chosen such that
  1. $\circ$
    $T_{k}\geq_{\text{a.s.}}T^{*}_{k}$, and
  2. $\circ$
Combining these statistics yields the estimator $\widehat{\nabla\Psi}(p)$ as defined in (2.8). Under the assumptions made above we have the following result regarding consistency of the gradient estimator.
Proposition 4.1.
Under Assumption 2.1, the estimator $\widehat{\nabla\Psi}(p)$ converges almost surely to $\nabla\Psi(p)$ as the interval size $T$ on which price $p$ is enforced goes to $\infty$.
The proof relies on applying a law of large numbers for regenerative processes and is deferred to Appendix A.4.
5 Convergence analysis of the learning algorithm
Proposition 5.1.
There exists constants $C_{1},C_{2}>0$ such that, for all $k\geq 1$,
$\nu_{k}$ $\leq C_{1},$ (5.3)
$|\beta_{k}|$ $\leq C_{2}\frac{\eta_{k-1}}{{T^{*}_{k-1}}^{2}}.$ (5.4)
Corollary 5.2.
5.1 Variability analysis
Therefore,
By definition,
Hence, by the continuous mapping theorem, for any $m\in{\mathbb{N}}$,
5.2 Bias analysis
In the analysis of the bias, the starting point is
[…]
$\leq\omega^{\rm(I)}_{k}+\overline{p}\,\omega^{\rm(II)}_{k}$, (5.6)
where
[…]
We aim to determine an upper bound on the rate at which $\beta_{k}$ vanishes as $k\to\infty$, which we do by analyzing the decay behavior of both $\omega_{k}^{\text{(I)}}$ and $\omega_{k}^{\text{(II)}}$.
5.2.1 Analysis of $\omega_{k}^{\rm(III)}$
In our analysis a key role is played by the following result.
Lemma 5.3.
As $k\to\infty$,
[…]
Proof.
As a first step, observe that $\omega_{1}^{\rm(III)}$ can be majorized via
[…]
where in the first inequality we have used Jensen’s inequality, and in the second inequality we have used the triangle inequality. In order to further dominate the expectation in (5.2.1), we construct a probability space $\chi$, and on this space we couple two M/G/1 + $H(p,V)$ queues with the same admission price $p_{k-1}$, interarrival-time seeds, and service times but with different initial workloads, as follows:
  1. $\circ$
  2. $\circ$
We now have a setup where Proposition 3.9 can be used. Specifically,
  1. $\circ$
  2. $\circ$
Substituting this into inequality (5.2.1), recalling that $\gamma_{1}<1$,
where in the last step we once more used the triangle inequality.
Now appealing to Proposition 3.9(b),
so that we arrive at the upper bound
$\omega_{k}^{\rm(III)}\leq\ \delta_{k-1}\leq\gamma_{1}^{N_{k-1}}\delta_{k-2}+\rho_{k-2},$ (5.9)
with, for any $i\in{\mathbb{N}}$, for ease suppressing the dependence on $k$,
$\delta_{i}$ $:=\operatorname{\mathbb{E}}\left[\left|W\left(\overline{T}_{i}\right)-\overline{W}_{\infty}(p_{i})\right|\ \big|\mathcal{F}_{k-1}\right],\quad\rho_{i}:=\operatorname{\mathbb{E}}\left[\big|\overline{W}_{\infty}(p_{i})-\overline{W}_{\infty}(p_{i+1})\big|\ \big|\mathcal{F}_{k-1}\right].$ (5.10)
By iteration we obtain that, the empty product being defined as $1$,
$\delta_{k-1}$ $\leq\gamma_{1}^{N_{1}+\cdots+N_{k-1}}\delta_{0}+\sum_{\ell=0}^{k-2}\rho_{\ell}\left(\prod_{m=\ell+2}^{k-1}N_{m}\right)=\mathcal{O}_{\operatorname{\mathbb{P}}}\left(\rho_{k-2}\right);$ (5.11)
$\rho_{k-2}\leq\mathcal{O}\left(\eta_{k-1}\frac{1}{{T^{*}_{k-1}}^{2}}\right).$ (5.12)
Upon combining the above, we conclude that the stated holds. ∎
5.2.2 Analysis of $\omega_{k}^{\rm(I)}$
Our next goal is to use Lemma 5.3 to obtain an upper bound on the rate of convergence of $\omega_{k}^{\rm(I)}$ as $k\to\infty$.
Lemma 5.4.
As $k\to\infty$,
Proof.
We now proceed to analyze both terms separately.
Then,
where we have used $\xi>\epsilon$ in the first inequality, and Lemma 5.3 in the final step.
we write
thus leading to the inequality
so that
Upon combining the above, we have arrived at the upper bound
5.2.3 Analysis of $\omega_{k}^{\rm(II)}$
We proceed by proving the following lemma.
Lemma 5.5.
As $k\to\infty$,
Proof.
where, with ${{\bm{1}}}_{k}$ as defined in the proof of Lemma 5.4,
so that
Recalling that $\xi_{2}>\epsilon$, and $\xi_{1}\leq G_{\max}$, we find
Therefore,
Lemma 5.6.
As $k\to\infty,$
6 Regret Analysis
We can decompose the above as
$R(L)=R_{1}1(L)+R_{2}(L)\ =\sum_{k=1}^{L}R_{1,k}+\sum_{k=1}^{L}R_{2,k},$
where
6.1 Analysis for $R_{1}(L)$
The following lemma establishes an upper bound on $R_{1}(L)$. Define
Lemma 6.1.
For any $L$,
Proof.
By the mean-value theorem for random variables [26], for a random variable $\xi_{k-1}$ taking values in $(0,1)$, and for $\widetilde{p}_{k-1}=\xi_{k-1}p^{*}+\big(1-\xi_{k-1}\big)p_{k-1}$,
$R_{1,k}$ $=T^{*}_{k}\operatorname{\mathbb{E}}\big[\Psi(p^{*})-\Psi(p_{k-1})\big]=T^{*}_{k}\operatorname{\mathbb{E}}\Big[(p^{*}-p_{k-1})\operatorname{\mathbb{E}}\big[\nabla\Psi\big(\widetilde{p}_{k-1}\big)\big]\Big]\leq T^{*}_{k}\operatorname{\mathbb{E}}\Big[\big|p^{*}-p_{k-1}\big|\operatorname{\mathbb{E}}\big|\nabla\Psi\big(\widetilde{p}_{k-1}\big)\big|\Big].$
With reference to Corollary 3.6,
$\frac{1}{\operatorname{\mathbb{E}}[A_{\infty}(\overline{p})]}-\overline{p}\ G_{\max}\frac{1}{\operatorname{\mathbb{E}}[A_{\infty}(\underline{p})]^{2}}\leq\nabla\Psi(p)$ $=\frac{1}{\operatorname{\mathbb{E}}[A_{\infty}(p)]}-p\frac{\nabla\operatorname{\mathbb{E}}\big[A_{\infty}(p)\big]}{\operatorname{\mathbb{E}}[A_{\infty}(p)]^{2}}\leq\frac{1}{\operatorname{\mathbb{E}}\big[A_{\infty}(\underline{p})\big]}.$
Therefore,
$R_{1,k}$ $\leq\max\left\{\frac{1}{\operatorname{\mathbb{E}}\big[A_{\infty}(\underline{p})\big]},\left|\frac{1}{\operatorname{\mathbb{E}}[A_{\infty}(\overline{p})]}-\overline{p}\ G_{\max}\frac{1}{\operatorname{\mathbb{E}}[A_{\infty}(\underline{p})]^{2}}\right|\right\}T^{*}_{k}\operatorname{\mathbb{E}}\big|p^{*}-p_{k-1}\big|$
$=\sigma\ T^{*}_{k}\operatorname{\mathbb{E}}\big|p^{*}-p_{k-1}\big|.$
From this, the result follows. ∎
6.2 Analysis for $R_{2}(L)$
In this subsection we state and prove an upper bound on $R_{2}(L)$.
Lemma 6.2.
For any $L$,
Proof.
In order to dominate the expectation above, we construct a probability space $\big(\Omega,\operatorname{\mathscr{G}}_{k-1},\operatorname{\mathbb{P}}\big)$ and on it, we couple two M/G/1  +  $H(p,V)$ queues with the same admission price $p_{k-1}$, but with different initial workloads.
  1. 1.
    The first queue with initial workload $W_{\infty}(p_{k-1})$ and counting process $N_{1}(\cdot)$ of effective arrivals.
  2. 2.
    The second queue with initial workload $W\left(\overline{T}_{k-1}\right)$ and counting process $N_{2}(\cdot)$ of effective arrivals.
Then,
$R_{2,k}$ $=\operatorname{\mathbb{E}}\Big[\Psi(p_{k-1})\Big]T^{*}_{k}-\operatorname{\mathbb{E}}\Big[p_{k-1}\operatorname{\mathbb{E}}\big[N\big(\overline{T}_{k-1},\overline{T}_{k-1}+T^{*}_{k}\big]\big|\operatorname{\mathscr{F}}_{k-1}\big]\Big]$
$=\operatorname{\mathbb{E}}\Big[p_{k-1}\operatorname{\mathbb{E}}\big[N_{1}\big(\overline{T}_{k-1},\overline{T}_{k-1}+T^{*}_{k}\big]-N_{2}\big(\overline{T}_{k-1},\overline{T}_{k-1}+T^{*}_{k}\big]\big|\operatorname{\mathscr{G}}_{k-1}\big]\Big]$
$\leq\overline{p}\ \operatorname{\mathbb{E}}\Big[\operatorname{\mathbb{E}}\Big[\left|N_{1}\big(\overline{T}_{k-1},\overline{T}_{k-1}+T^{*}_{k}\big]-N_{2}\big(\overline{T}_{k-1},\overline{T}_{k-1}+T^{*}_{k}\big]\right|\ \Big|\operatorname{\mathscr{G}}_{k-1}\Big]\Big].$
From Proposition 3.9(d), we know that
$\operatorname{\mathbb{E}}\Big[\big|N_{1}\big(\overline{T}_{k-1},\overline{T}_{k-1}+T^{*}_{k}\big]-N_{2}\big(\overline{T}_{k-1},\overline{T}_{k-1}+T^{*}_{k}\big]\big|\ \Big|\operatorname{\mathscr{G}}_{k-1}\Big]\leq\frac{\Lambda H\big(\underline{p},0\big)}{1-\gamma_{1}}\operatorname{\mathbb{E}}\left|W_{\infty}\left(p_{k-1}\right)-W\left(\overline{T}_{k-1}\right)\right|,$
from which we can conclude that
$R_{2,k}\leq\overline{p}\ \frac{\Lambda H\big(\underline{p},0\big)}{1-\gamma_{1}}\operatorname{\mathbb{E}}\left|W_{\infty}\left(p_{k-1}\right)-W\left(\overline{T}_{k-1}\right)\right|.$
6.3 Analysis for $R(L)$
Assumption 6.3.
  1. (a)
    $\left(1+\frac{1}{k}\right)^{\alpha}\leq 1+\frac{K_{0}}{2}\eta_{k}$,
  2. (b)
    $B_{k}\leq\frac{K_{0}}{8}k^{-\alpha}$.
Theorem 6.4.
Under Assumption 6.3, for any $k\geq 1$,
Proof.
Deferred to Appendix A.5. ∎
Theorem 6.4 gives us the rate of convergence of iterates $\{p_{k}\}_{k\geq 1}$ to $p^{*}$. This rate is determined by $\alpha$ which comes from Assumption 6.3. By Equation (5.4), choosing a sequence $\big\{T^{*}_{k}\big\}_{k\geq 1}$ which grows fast tightens the rate at which the bias decays. This allows a potentially larger value of $\alpha$ to satisfy Assumption 6.3(b), and thereby to obtain a faster convergence rate to $p^{*}$. What this means in words is - when the chosen window sizes are large enough, the gradient estimators $\widehat{\nabla\Psi}(p)$ are more accurate allowing for faster convergence to $p^{*}$. Recall that the service operator’s goal is to minimize the regret which is accomplished by learning the optimal price as quickly as possible. Clearly then, there is a trade-off to study here: large window sizes take fewer iterations to learn the optimal price. However, it could potentially take more time, and we may therefore be favouring smaller window sizes more. We explore this tradeoff in Section 7.3.
Theorem 6.5.
Under Assumption 6.3, as $L\rightarrow\infty$,
$R(L)=\mathcal{O}\left(\sum_{k=1}^{L}T^{*}_{k}\ k^{-\alpha/2}\right).$
Proof.
Using Theorem 6.4 and Jensen’s inequality,
which implies, in conjunction with Lemma 6.1 that, as $L\to\infty$,
$R_{1}(L)=\mathcal{O}\left(\sum_{k=1}^{L}T^{*}_{k}\ k^{-\alpha/2}\right).$
Equations (5.10), (5.11), and (5.12), together with Lemma 6.2, lead to
as $L\to\infty$. Upon combining the above two displays,
as $L\to\infty$. ∎
7 Numerical experiments
In the preceding sections, we provided a comprehensive mathematical analysis of our stochastic optimization algorithm. Section 7.1 empirically investigates how the service-time distribution influences the optimal price $p^{*}$. In Section 7.2, we study the algorithm’s performance through a series of illustrative examples that demonstrate its convergence behavior. Subsequently, in Section 7.3, we investigate the empirical impact of different window size choices on the convergence rate toward the optimal price, offering insights into how this parameter influences algorithmic efficiency. We investigate two examples of joining functions: the exponential form of Example A, and the following polynomial form.
Example B.
Remark 7.1.
This section includes several examples based on Gamma-distributed service times. We emphasize that our notation $\mathrm{Gamma}\bigl(\alpha,\beta\bigr)$ denotes a Gamma distribution with shape parameter $\alpha$ and rate parameter $\beta$. Its mean is $\alpha/\beta$, and its variance is $\alpha/\beta^{2}$. We recall that ${\rm Exp}(\beta)$ denotes an exponentially distributed random variable with mean $1/\beta$ (and hence variance $1/\beta^{2}$).
7.1 Impact of service-time distribution
In this subsection, our goal is to understand how the service-time distribution $G(\cdot)$ impacts the optimal price $p^{*}$ for the system. To this end, Figure 6 presents three sets of plots of the objective function $\Psi(p)$, each corresponding to a different family of service-time distributions.
To the best of our knowledge, no closed-form expression for $\Psi(p)$ is available, not even for the simplest setting of Example A. Therefore, to obtain a reliable approximation of $\Psi(p)$ and thereby, the optimal price $p^{*}$, we simulate each example over a fine enough grid of prices for $N_{\text{eff}}=5\times 10^{5}$ effective arrivals each. We then compute the empirical average revenue per unit time $\widehat{\Psi}(p):=p/\widehat{A_{\infty}}(p)$, where $\widehat{A_{\infty}}(p)$ is the sample average of the $N_{\text{eff}}$ effective interarrival times. Finally, $p^{*}$ is approximated from the grid as the price which maximizes $\widehat{\Psi}(p)$. Throughout these experiments, all other model parameters are held fixed. In particular, we choose $\Lambda=20$ and $H(p,V)=\exp\bigl(-0.1p-0.2V\bigr)$.
  1. 1.
    $p^{*}$ increases whenever the mean service time increases.
  2. 2.
  3. 3.
    $\Psi(p)$ (and in particular $\Psi(p^{*})$) both increase whenever the mean service time increases.
  4. 4.
    $\Psi(p)$ (and in particular $\Psi(p^{*})$) both decreases whenever the variance of the service time increases.
7.2 Convergence of price iterates
Feature Example 1 Example 2 Example 3 Example 4
$\Lambda$ 2020 2020 2020 2020
$N_{\text{eff}}$ $5\times 10^{5}$ $5\times 10^{5}$ $5\times 10^{5}$ $5\times 10^{5}$
$p^{*}$ $\approx 9.3$ $\approx 29.0$ $\approx 29.5$ $\approx 16.5$
$\Psi(p^{*})$ $\approx 16.8$ $\approx 17.2$ $\approx 17.8$ $\approx 9.4$
$T^{*}_{k}$ $50\log(k+1)$ $50\log(k+1)$ $50\log(k+1)$ $50\log(k+1)$
$\eta_{k}$ $20/k^{0.75}$ $20/k^{0.75}$ $20/k^{0.75}$ $20/k^{0.75}$
Table 1: Model specifications for Examples 1–4.
Remark 7.2.
Certain choices of the service time distribution $G(\cdot)$ and the joining probability $H\big(p,V\big)$ give rise to plots for $\Psi(p)$ which plateau around $p^{*}$. In such cases, the algorithm naturally has more difficulty converging precisely to $p^{*}$. Example 4 illustrates this phenomenon. Figure 10(c) shows $\widehat{\Psi}(p)$ in a neighborhood of $p^{*}$, where the objective function is visibly flat. Consequently, as seen in Figure 10(b), the algorithm does not converge fully to $p^{*}$ even after 500 iterations, in contrast to the other examples in which convergence is much faster. However, it is important to note that, from a practical standpoint, this behavior is not problematic: the service operator’s primary goal is to maximize revenue. In that respect, the algorithm performs well, quickly moving into the vicinity of the optimal price, where the resulting revenue is essentially optimal.
(a) $\widehat{\Psi}(p)$ versus pp.
(b) $p_{k-1}$ vs. $k$ for $L=150$ iterations, $p_{0}=20$.
Figure 7: SGD-based learning algorithm in action for Example 1.
(a) $\widehat{\Psi}(p)$ versus pp.
(b) $p_{k-1}$ vs. $k$ for $L=100$ iterations, $p_{0}=50$.
Figure 8: SGD-based learning algorithm in action for Example 2.
(a) $\widehat{\Psi}(p)$ versus pp.
(b) $p_{k-1}$ vs. $k$ for $L=100$ iterations, $p_{0}=10$.
Figure 9: SGD-based learning algorithm in action for Example 3.
(a) $\widehat{\Psi}(p)$ versus pp.
(b) $p_{k-1}$ vs. $k$ for $L=500$ iterations, $p_{0}=20$.
(c) $\widehat{\Psi}(p)$ versus $p$ in the neighbourhood of $p^{*}$.
Figure 10: SGD-based learning algorithm in action for Example 4.
7.3 Convergence time comparison
  • $\circ$
    Let us first consider Figure 11(a). In Section 7.2, our attention was on plots of the price iterates as functions of the iteration number. In contrast, to better understand how quickly the iterates converge to $p^{*}$ in time, Figure 11(a) presents the price iterates against the actual time. The figure displays the trajectories produced by the window choices $T^{*}_{k}=10\sqrt{k}$, $T^{*}_{k}=50\sqrt{k}$, and $T^{*}_{k}=100\sqrt{k}$, with the coloured bands indicating the dispersion of the corresponding iterates. A wider band indicates more fluctuation, while a narrower band indicates stability.
    The figure contains a second plot showing the evolution of the error $\lvert p_{k-1}-p^{*}\rvert$ as a function of time. From these results, it is apparent that $C=10$ (blue curve) yields the fastest convergence, followed by the $C=50$ (red curve) and $C=100$ (green curve) converging the slowest. An analogous pattern is observed for the logarithmic window-sizes in Figure 11(b).
    (a) Price iterates versus simulation time for $T^{*}_{k}=C\sqrt{k}$, $p_{0}=30$
  • $\circ$
    Varying window size rates. Next, we vary the rate at which the window sizes grow by studying the behavior of the price iterates over time for window functions of the form $T^{*}_{k}=C\,k^{\alpha}$, considering multiple values of the exponent $\alpha$. The corresponding plots are shown in Figure 12. Once again, similar to the observation in Figure 12, we observe that $\alpha=0.1$ (blue curve) yields the fastest convergence followed by $\alpha=0.3$ (red curve) and $\alpha=0.5$ being the slowest.
The above two experiments highlight an inherent trade-off: increasing $C$ and $e$ leads to longer time windows, which provide more data per iteration and thus yield more accurate gradient estimates. However, this comes at the cost of fewer updates within a fixed time budget. From Figures 1112, it becomes evident that smaller values of $C$ and $e$ — i.e., slower growth of the window size — lead to faster convergence of the price iterates to the optimal price.
8 Concluding remarks
  • $\circ$
  • $\circ$
  • $\circ$
— APPENDICES —
Appendix A Proofs
A.1 Proof of Proposition 2.2
Proof of Proposition 2.2.
Therefore, $\left(\overline{W}_{i},A_{i+1}\right)$ is an ergodic sequence. Hence, the ergodic theorem entails that
Then Equation (2.5) follows due to
This proves the stated. ∎
A.2 Proof associated with Example A
Then, as per the definition of $\Xi[\zeta]$ in Example A,
A.3 Proof of Lemma 3.5
Proof of Lemma 3.5.
First, we show via induction that for all $n\geq 0$,
Denoting
we can rewrite
which is an increasing linear map. As a result,
This completes the proof. ∎
A.4 Proof of Proposition 4.1
@@ A.2 Proof associated with Example A
A.5 Proof of Theorem 6.4
First, we show that there exists $\eta<\infty$ such that hypothesis (a) of Theorem 6.4 is satisfied. Consider the function

### v2

@@ A.2 Proof corresponding to Example A
    $<\left(\overline{W}^{1}_{j}-\overline{W}^{2}_{j}\right).$ (A.1)
@@
    $N_{2}\left(0,t\right]-N_{1}\left(0,t\right]=N_{2}\left(t-\Delta,t\right]\leq_{\text{st}}\overline{N}\left(\Delta\right),$ (A.2)
    where the random variable $\overline{N}(\Delta)$ records the number of arrivals in an interval of length $\Delta$, under a Poisson stream of arrivals with rate $\Lambda H\big(\underline{p},0\big)$. The stochastic dominance follows from an evident coupling argument: each interarrival time in the second queue stochastically dominates an $\text{Exp}\big(\Lambda\,H\big(\underline{p},0\big)\big)$ random variable. Similarly, under the complementary event $\big\{\overline{W}^{1}_{0}<\overline{W}^{2}_{0}\big\}$ we have (A.2) as well. In conclusion, $\big|N_{1}\left(0,T^{*}\right]-N_{2}\left(0,T^{*}\right]\big|\leq_{\text{st}}\overline{N}\left(\Delta\right).$
@@
A.5 Proof of Proposition 4.2
@@
A.6 Proofs for Section 5.2 - Bias Analysis
Recall that the upper bound on the bias could be decomposed into the following terms,
[…]
Before establishing the convergence rates of these terms, as stated in Lemmas 5.3 and 5.4 we define
[…]
which we analyze fist as an intermediate step.
Lemma A.1.
As $k\to\infty$,
$\omega_{k}^{\rm(III)}=\mathcal{O}\left(\eta_{k-1}\frac{1}{{T^{*}_{k-1}}^{2}}\right).$ (A.3)
Proof.
As a first step, observe that $\omega_{1}^{\rm(III)}$ can be majorized via
[…]
where in the first inequality we have used Jensen’s inequality, and in the second inequality we have used the triangle inequality. In order to further dominate the expectation in (A.6), we construct a probability space $\chi$, and on this space we couple two M/G/1 + $H(p,V)$ queues with the same admission price $p_{k-1}$, interarrival-time seeds, and service times but with different initial workloads, as follows:
  1. $\circ$
  2. $\circ$
We now have a setup where Proposition 3.9 can be used. Specifically,
  1. $\circ$
  2. $\circ$
Substituting this into inequality (A.6), recalling that $\gamma_{1}<1$,
where in the last step we once more used the triangle inequality.
Now appealing to Proposition 3.9(b),
so that we arrive at the upper bound
$\omega_{k}^{\rm(III)}\leq\ \delta_{k-1}\leq\gamma_{1}^{N_{k-1}}\delta_{k-2}+\rho_{k-2},$ (A.5)
with, for any $i\in{\mathbb{N}}$, for ease suppressing the dependence on $k$,
$\delta_{i}$ $:=\operatorname{\mathbb{E}}\left[\left|W\left(\overline{T}_{i}\right)-\overline{W}_{\infty}(p_{i})\right|\ \big|\mathcal{F}_{k-1}\right],\quad\rho_{i}:=\operatorname{\mathbb{E}}\left[\big|\overline{W}_{\infty}(p_{i})-\overline{W}_{\infty}(p_{i+1})\big|\ \big|\mathcal{F}_{k-1}\right].$ (A.6)
By iteration we obtain that, the empty product being defined as $1$,
$\delta_{k-1}$ $\leq\gamma_{1}^{N_{1}+\cdots+N_{k-1}}\delta_{0}+\sum_{\ell=0}^{k-2}\rho_{\ell}\left(\prod_{m=\ell+2}^{k-1}N_{m}\right)=\mathcal{O}_{\operatorname{\mathbb{P}}}\left(\rho_{k-2}\right);$ (A.7)
$\rho_{k-2}\leq\mathcal{O}\left(\eta_{k-1}\frac{1}{{T^{*}_{k-1}}^{2}}\right).$ (A.8)
The proof is concluded by combining (A.5), (A.7) and (A.8). ∎
Proof of Lemma 5.3.
We now proceed to analyze both terms separately.
Then,
where we have used $\xi>\epsilon$ in the first inequality, and Lemma A.1 in the final step.
we write
thus leading to the inequality
so that
Upon combining the above, we have arrived at the upper bound
Proof of Lemma 5.4.
where, with ${{\bm{1}}}_{k}$ as defined in the proof of Lemma 5.3,
so that
Recalling that $\xi_{2}>\epsilon$, and $\xi_{1}\leq G_{\max}$, we find
Therefore,
A.7 Proof of Theorem 6.4
First, we show that there exists $\eta<\infty$ such that Assumption 6.3(a) is satisfied. Consider the function